OpenStudy (anonymous):

how do you simplify x^8/5*x^4/5 / x^2/5?

OpenStudy (anonymous):

what is \(\frac{8}{5}+\frac{4}{5}\) ?

OpenStudy (anonymous):

can't be too bad since the denominators are the same

OpenStudy (anonymous):

\[\frac{ 12 }{ 5 }\] or \[\frac{ 12 }{ 10 }\]

OpenStudy (anonymous):

no, not "or" it is \(\large \frac{12}{5}\) next step, subtract \(\large \frac{2}{5}\)